i am really stumped here!
f(x)=e^2x
what is f(1/4)?

Answer by rapaljer(4551)

If f(x)=e^2x, then
f(1/4) just means to substitute the value of 1/4 in place of the x.
f(1/4)=e^(2*1/4)
f(1/4)=e^(1/2)=e^0.5=1.6487 approximately!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus